### v2

@@ 4.1 Notation and problem statement
We denote the full state of the atmosphere at a particular time $t$ as a tensor $\mathbf{x}(t)$ with dimensions $V\times H\times W$, where $V$ represents the total number of variables, $H$ indicates the number of latitude coordinates, and $W$ signifies the number of longitude coordinates. The indexing scheme $\mathbf{x}_{v,i,j}(t)$ refers to the state of variable $v$ at time $t$ and the latitude-longitude coordinates $(i,j)$. Additionally, we represent GDAS prepbufr conventional atmospheric observations at time $t$ as a tensor $\mathbf{y}^{c}(t)$ of the same dimensions $V\times H\times W$. For unobserved variable $v$ at coordinates $(i,j)$, we set $\mathbf{y}^{c}_{v,i,j}(t)=\text{NAN}$, while for observed variable $v$, $\mathbf{y}^{c}_{v,i,j}(t)=\mathbf{x}_{v,i,j}(t)+\mathbf{\varepsilon}_{v,i,j}$, where $\mathbf{\varepsilon}_{v,i,j}$ represents the observation error. For satellite observations, we denote this as $\mathbf{y}^{s}(t)=\mathcal{H}^{s}\left(\mathbf{x}(t)\right)+\mathbf{\varepsilon}^{s}$, where $\mathbf{\varepsilon}^{s}$ is the satellite observation error and $\mathcal{H}^{s}$ is the observation operator that maps atmospheric variables to satellite observation variables.
We aim to develop a fully AI-driven global weather forecasting system that integrates both forecasting and DA models. For the forecasting model, given a system state $\mathbf{x}(t)$, our objective is to predict the state $\mathbf{x}(t+\Delta t)$ at a future time $t+\Delta t$. To accomplish this, the forecasting model $\mathcal{M}_{t\rightarrow t+\Delta t}:\mathbf{x}(t)\mapsto\hat{\mathbf{x}}(t+\Delta t)$ is trained to map the state of the atmosphere at present to a predicted state $\hat{\mathbf{x}}(t+\Delta t)$ for the future time $t+\Delta t$. Regarding the assimilation model, given a background field $\mathbf{x}^{b}(t_{0})$ at time $t_{0}$, GDAS prepbufr observations $\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\}$, and satellite observations $\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}$ during the DAW $t_{0},\cdots,t_{K}]$, our goal is to learn to estimate the optimal initial field $\mathbf{x}^{a}(t_{0})$ by fusing these data. To assimilate the satellite observations, we need to learn an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. Subsequently, we need to learn a DA model $\mathcal{D}:\left(\mathbf{x}^{b}(t_{0}),\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\},\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}\right)\mapsto\mathbf{x}^{a}(t_{0})$, where the reanalysis serves as the reference.
@@ 4.2 Dataset
For the AMSU-A and MHS observations, nearest neighbor interpolation was utilized to map both the raw satellite data and the auxiliary information onto a 1.40625 spatial resolution and 3-hour temporal resolution interval. In this study, we excluded observations from regions located above 60 north and south latitude to avoid complications due to sea ice  [noh2023assimilation]. Additionally, we implemented a straightforward quality control procedure that eliminated data with brightness temperature values exceeding 350 K or falling below 150 K. Since MHS observations are highly influenced by clouds, a basic cloud-detection method was applied to exclude cloud-contaminated portions of the MHS data. For details on the cloud-detection methodology, please refer to the reference paper [ma2013introduction]. For the 10-meter wind field data obtained from ASCAT and the upper-air wind field data obtained from SATWIND observations, we adopt an identical processing methodology to that implemented for GDAS prepbufr observational datasets.
@@
All models are trained on data collected until 2021, with the year 2022 designated as a validation set. Subsequently, data from 2023 is utilized for testing. For details of the reanalysis dataset as well as observations used in this study, please see Supplementary Text and Tables S1 to S4.
The TC ground truth used in this study is obtained from the IBTrACS dataset, accessible at here, which compiles recent and historical TC records from multiple agencies and provides the most comprehensive global archive of TC information.
4.3 The fully AI-driven global weather forecasting system
The XiChen model is based on our proposed Conditional Hybrid Neural Operator (CHNO), which incorporates cross-attention to integrate conditional information into our Hybrid Neural Operator (HNO). This design enables a unified model architecture adaptable to multiple downstream tasks. Additionally, since the Adaptive Fourier Neural Operator (AFNO) [guibas2021adaptive] performs only global feature extraction through convolution operations in the Fourier space, we introduce a convolution-based branch to enhance local feature extraction capabilities, thereby developing the HNO. For further details on the CHNO and XiChen model architectures, refer to Figure S2. A comparison between HNO and AFNO in the context of medium-range weather forecasting is provided in the Supplementary Text and Figure S4.
Figure 9: Overview of our XiChen. (a) The overall framework of the fully AI-driven global weather forecasting system “XiChen”. XiChen integrates background field data, conventional observations, and indirect satellite observations within the DAW to perform the DA task. The process begins by computing the 4DVar cost function, and it subsequently determines the gradient with respect to the background field. By combining this gradient with the background field, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. The green box denotes the weather forecasting model, the purple represents the observational operator, and the blue signifies the DA model. XiChen undergoes three stages of training: (b) It is pre-trained as a medium-range weather forecasting model, with lead time ($\Delta t$) as the conditional input; (c) It is fine-tuned as an observation operator, using auxiliary observational information of the satellite as the conditional input; (d) It is fine-tuned as a DA model, with the gradient of the 4DVar cost function as the input and the background field as the conditional input; (e) We provide a dual DA framework to operationalize XiChen as a cycling forecasting system.
Our fully AI-driven global weather forecasting system uses XiChen to cover core components of the NWP pipeline, from observation operators and DA models to the weather forecasting model. As shown in Figure 9(a), XiChen integrates background field data, conventional observations, and satellite observations within the DAW to perform the DA task. The process begins by computing the 4DVar cost function, and it subsequently determines the gradient with respect to the background field. Here, the 4DVar cost function $\mathcal{J}^{\text{4DVar}}(\mathbf{x}(t_{0}))$ represents as follows:
$\tiny{\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))=\frac{1}{2}\|\mathbf{x}^{b}(t_{0})-\mathbf{x}(t_{0})\|^{2}_{\mathbf{B}^{-1}}+\frac{1}{2}\sum^{K}_{k=0}\|\mathbf{y}(t_{k})-\mathcal{H}\left(\mathcal{M}_{t_{0}\rightarrow t_{k}}(\mathbf{x}(t_{0}))\right)\|^{2}_{\mathbf{R}^{-1}},}$ (1)
where $\mathbf{x}(t_{0})$ denotes the initial field to be optimized, $\mathbf{x}^{b}(t_{0})$ represents the background field, and $\mathbf{y}(t_{k})$ represents the observations at time $t_{k}$. The forecasting model, denoted as $\mathcal{M}_{t_{0}\rightarrow t_{k}}$, maps the initial field to the field at time $t_{k}$. Additionally, $\mathbf{B}$ and $\mathbf{R}$ represent the background and observation error covariance matrices, respectively. By combining the gradient $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ with the background field $\mathbf{x}^{b}(t_{0})$, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. Among them, our DA task involves sequentially employing DA models for different observational data. For example, in this study, XiChen first assimilates GDAS prepbufr data, then assimilates MHS, ASCAT, AMSU-A, and SATWND observations. This configuration enhances XiChen’s scalability. The assimilation of new observations in the future merely requires fine-tuning the respective observation operators and DA model components, eliminating the need to retrain the entire DA system. For details of the cascade DA framework, please refer to the Supplementary Text and Figure S1 (b).
The following subsections will provide a detailed explanation of the medium-range weather forecast pre-training, observation operator fine-tuning, and DA model fine-tuning.
@@ 4.3 The fully AI-driven global weather forecasting system
4.4 Pre-training for Medium-range Weather Forecasting
@@
We aim to design a foundation model that can be pre-trained on the reanalysis data and then be finetuned to solve various downstream weather tasks to build a fully AI-driven global weather forecasting system. As illustrated in Figure 9(b), during the pre-training process, the foundation model takes the reanalysis field $\mathbf{x}(t)$ as the input and the lead time $\Delta t$ as the condition and outputs the predicted state $\hat{\mathbf{x}}(t+\Delta t)$. In this study, we sample $\Delta t$ from the set $\{1,3,6,12,24\}$ hours. This process enables the foundation model to effectively capture the relationships among various weather variables while accounting for both the dynamical and thermodynamic characteristics of the atmosphere. To generate predictions for any time increments, the predictions from the learned forecasting model can be combined using the hierarchical temporal aggregation methodology proposed by Pangu-Weather [bi2023accurate]. Thus, the foundation model is pre-trained to minimize the following loss:
@@
$\tiny{\mathcal{L}(\theta)=\mathbb{E}\left[\frac{1}{VHW}\sum^{V}_{v=1}\sum^{H}_{h=1}\sum^{W}_{w=1}\omega(v)L(i)\|\Delta\mathbf{x}_{v,i,j}(t+\Delta t)\|_{1}\right],}$ (2)
@@
where $\omega(v)$ denotes the pressure-weighted of variable $v$ [lam2023learning], $\Delta\mathbf{x}(t+\Delta t)=\mathcal{M}_{t\rightarrow t+\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\Delta t)$ is the difference between the predicted state and the ground truth at time $t+\Delta t$, $\|\|_{1}$ represents the $L_{1}$-norm, and $L(i)$ is the latitude-weighting factor [rasp2020weatherbench]:
@@
where $\text{lat}(i)$ is the latitude of the $i$th row of the grid. This term is commonly used in the training process of previous AI-based weather forecasting models [kurth2023fourcastnet, bi2023accurate, lam2023learning]. This pre-training process costs about 45 hours to run on 8 A100-40GB GPUs.
@@
where $\tau$ denotes the roll out step and $\Delta\mathbf{x}_{v,i,j}(t+\tau\Delta t)=\mathcal{M}_{t\rightarrow t+\tau\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\tau\Delta t)$.
In practice, we first fine-tune the pre-trained forecasting model with roll-out steps from $T=2$ to $T=4$ to develop the XiChen-Short model. The weights obtained from the XiChen-Short model are then used to initialize the XiChen-Medium model, which is subsequently fine-tuned to achieve optimal forecasting performance for periods of 5 to 10 roll-out steps. We use the same sampled value of $\Delta t$ for all $T$ steps. Our forecasting model is similar to FuXi [fudanchen2023fuxi] in its forecasting process. Since the analysis field errors provided by the XiChen DA model are comparable to those of XiChen-Short’s 2-day forecasts initialized with ERA5, the 10-day medium-range weather forecasts using initial fields estimated by the XiChen DA model are generated by XiChen-Short for the first 3 days and by XiChen-Medium for the remaining 7 days. The fine-tuning process of the XiChen forecasting model costs about 94 hours to run on 8 A100-40GB GPUs. For additional information on the model training configuration, please refer to the Supplementary Text.
@@ 4.5 Fine-tuning for the Observation Operator
To compare the atmospheric state with satellite observations, the observation operator is essential [liang2023machine]. The simulated radiance is computed at each observation location using a profile of atmospheric variables. We fine-tune the pre-trained model by taking the satellite’s attitude parameters and scan position as conditions to construct an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. The loss function used to train $\hat{\mathcal{H}}^{s}$ is as follows:
@@ 4.6 Fine-tuning for Data Assimilation
Notably, we leverage the atmospheric dynamics information provided by the XiChen forecast model by incorporating 4D-Var a priori knowledge. This approach effectively propagates observations across the atmospheric field’s spatial and variable dimensions. Furthermore, the 4D-Var cost function enables the seamless integration of diverse observational operators, facilitating the assimilation of raw satellite observations.
As illustrated in Figure 9(c), we fine-tune the pre-trained model to function as a DA model by incorporating the gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$. This approach follows the derivation presented in 4DVarFormer [wang2024accurate]. However, unlike 4DVarFormer, XiChen outputs the representation of analysis increments in latent space rather than the analysis increments themselves. This representation is added to the embedding of the background field $\mathbf{x}^{b}(t_{0})$ and subsequently decoded by the HNO blocks to generate the analysis field.
@@
4.7 Dual DA framework to operationalize XiChen as a cycling forecasting system
As shown in Figure 9(e), to circumvent the temporal latency inherent in AI-based 4DVar methods [xiao2024fengwu4dvar, wang2024accurate] when deploying XiChen as a continuous forecasting system for acquiring medium-range prognostic results, we have architected a dual assimilation framework. Within this framework, we establish a configuration encompassing a 12-hour DAW and a 6-hour DAW. In the 12-hour DAW, we systematically assimilate observations retrospectively from the analysis time point at intervals of 0, 3, 6, and 9 hours to derive an accurate analysis field. This analysis field serves as the input for XiChen to generate a 12-hour forecast, subsequently yielding the background field for the initialization of the subsequent DAW cycle. Concurrently, within the 6-hour DAW configuration, we assimilate observations retrospectively from the analysis time point at intervals of 0 and 3 hours to generate the initial condition to initialize XiChen for the provision of medium-range forecast products. The 12-hour DAW employs 11 ensemble members, each derived from the previous cycle’s analysis field with added perturbations, processed through a 12-hour forecast. The perturbation method, similar to that used in Pangu [bi2023accurate], generates perturbations using 3 octaves of Perlin noise with scales of 0.2, 0.1, and 0.05, and the number of periods along each axis (longitude and latitude) is set to 1, 2, and 4, respectively. The weight of each member is calculated as follows:
$\mathcal{J}^{\text{3DVar}}_{O}(\mathbf{x}(t_{0}))$ $=\frac{1}{2}\|\mathbf{y}(t_{0})-\mathcal{H}\left(\mathbf{x}(t_{0})\right)\|^{2}_{\mathbf{R}^{-1}},$ (7)
$\omega_{idx}$ $=\frac{\exp(-J^{3DVar}_{O}(\mathbf{x}_{idx}(t_{0})))}{\sum^{Ens}_{idx^{\prime}=0}\exp(-J^{3DVar}_{O}(\mathbf{x}_{idx^{\prime}}(t_{0})))},$ (8)
where $\mathcal{J}^{\text{3DVar}}_{O}(\mathbf{x}(t_{0}))$ represents the observation term of the three-dimensional variational cost function, $Ens$ is the number of ensemble members, idxidx and $idx^{\prime}$ are the indices of the ensemble members, $\omega_{idx}$ represents the weight of the idxidx ensemble member. In this context, $\mathbf{y}(t_{0})$ and $\mathbf{x}(t_{0})$ are the GDAS prepbufr observations and the ensemble analysis member at the analysis time, respectively.
This dual DA framework facilitates the real-time acquisition of medium-range weather forecasting results without necessitating the temporal delay associated with the availability of all observations within the complete 12-hour DAW.

### v3

@@ 4.1 Notation and problem statement
We denote the full state of the atmosphere at a particular time $t$ as a tensor $\mathbf{x}(t)$ with dimensions $V\times H\times W$, where $V$ represents the total number of variables, $H$ indicates the number of latitude coordinates, and $W$ signifies the number of longitude coordinates. The indexing scheme $\mathbf{x}_{v,i,j}(t)$ refers to the state of variable $v$ at time $t$, and the latitude-longitude coordinates $(i,j)$. Additionally, we represent GDAS prepbufr conventional atmospheric observations at time $t$ as a tensor $\mathbf{y}^{c}(t)$ of the same dimensions $V\times H\times W$. For unobserved variable $v$ at coordinates $(i,j)$, we set $\mathbf{y}^{c}_{v,i,j}(t)=\text{NAN}$, while for observed variable $v$, $\mathbf{y}^{c}_{v,i,j}(t)=\mathbf{x}_{v,i,j}(t)+\mathbf{\varepsilon}_{v,i,j}$, where $\mathbf{\varepsilon}_{v,i,j}$ represents the observation error. For satellite observations, we denote this as $\mathbf{y}^{s}(t)=\mathcal{H}^{s}\left(\mathbf{x}(t)\right)+\mathbf{\varepsilon}^{s}$, where $\mathbf{\varepsilon}^{s}$ is the satellite observation error and $\mathcal{H}^{s}$ is the observation operator that maps atmospheric variables to satellite observation variables.
We aim to develop a global weather observation-to-forecast ML system that integrates both forecasting and DA models. For the forecasting model, given a system state $\mathbf{x}(t)$, our objective is to predict the state $\mathbf{x}(t+\Delta t)$ at a future time $t+\Delta t$. To accomplish this, the forecasting model $\mathcal{M}_{t\rightarrow t+\Delta t}:\mathbf{x}(t)\mapsto\hat{\mathbf{x}}(t+\Delta t)$ is trained to map the state of the atmosphere at present to a predicted state $\hat{\mathbf{x}}(t+\Delta t)$ for the future time $t+\Delta t$. Regarding the assimilation model, given a background field $\mathbf{x}^{b}(t_{0})$ at time $t_{0}$, GDAS prepbufr observations $\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\}$, and satellite observations $\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}$ during the DAW $t_{0},\cdots,t_{K}]$, our goal is to learn to estimate the optimal initial field $\mathbf{x}^{a}(t_{0})$ by fusing these data. To assimilate the satellite observations, we need to learn an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. Subsequently, we need to learn a DA model $\mathcal{D}:\left(\mathbf{x}^{b}(t_{0}),\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\},\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}\right)\mapsto\mathbf{x}^{a}(t_{0})$, where the reanalysis serves as the reference.
@@ 4.2 Dataset
For the AMSU-A and MHS observations, nearest neighbor interpolation was utilized to map both the raw satellite data and the auxiliary information onto a 1.40625 spatial resolution and 3-hour temporal resolution interval. In this study, we excluded observations from regions located above 60 north and south latitude to avoid complications due to sea ice  [noh2023assimilation]. Additionally, we implemented a straightforward quality control procedure that eliminated data with brightness temperature values exceeding 350 K or falling below 150 K. Since MHS observations are highly influenced by clouds, a basic cloud-detection method was applied to exclude cloud-contaminated portions of the MHS data. For details on the cloud-detection methodology, please refer to the reference paper [ma2013introduction]. For the 10-meter wind field data obtained from ASCAT and the upper-air wind field data obtained from SATWND observations, we adopt an identical processing methodology to that implemented for GDAS prepbufr observational datasets.
@@
All models are trained on data collected until 2021, with the year 2022 designated as a validation set. Subsequently, data from 2023 is utilized for testing. For details of the reanalysis dataset as well as observations used in this study, see Supplementary Text and Tables S1 to S4.
@@
4.3 The ML-based global weather assimilation and medium-range forecasting system
The XiChen model is based on our proposed Conditional Hybrid Neural Operator (CHNO), which incorporates cross-attention to integrate conditional information into our Hybrid Neural Operator (HNO). This design enables a unified model architecture adaptable to multiple downstream tasks. Additionally, since the Adaptive Fourier Neural Operator (AFNO) [guibas2021adaptive] performs only global feature extraction through convolution operations in the Fourier space, we introduce a convolution-based branch to enhance local feature extraction capabilities, thereby developing the HNO. For further details on the CHNO and XiChen model architectures, refer to Figure S2. A comparison between HNO and AFNO in the context of medium-range global weather forecasting is provided in the Supplementary Text and Figure S5.
Our system uses XiChen to cover core components of the NWP pipeline, from observation operators and DA models to the weather forecasting model. As shown in Figure 1(b), XiChen integrates background field data, conventional observations, and satellite observations within the DAW to perform the DA task. The process begins by computing the 4DVar cost function, which measures the misfit between observations and model forecasts. The 4DVar cost function $\mathcal{J}^{\text{4DVar}}(\mathbf{x}(t_{0}))$ is defined as follows:
$\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))=\frac{1}{2}\|\mathbf{x}^{b}(t_{0})-\mathbf{x}(t_{0})\|^{2}_{\mathbf{B}^{-1}}+\frac{1}{2}\sum^{K}_{k=0}\|\mathbf{y}(t_{k})-\mathcal{H}\left(\mathcal{M}_{t_{0}\rightarrow t_{k}}(\mathbf{x}(t_{0}))\right)\|^{2}_{\mathbf{R}^{-1}},$ (1)
where $\mathbf{x}(t_{0})$ denotes the initial field to be optimized, $\mathbf{x}^{b}(t_{0})$ represents the background field, and $\mathbf{y}(t_{k})$ represents the observations at time $t_{k}$. The forecasting model, denoted as $\mathcal{M}_{t_{0}\rightarrow t_{k}}$, maps the initial field to the field at time $t_{k}$. Additionally, $\mathbf{B}$ and $\mathbf{R}$ represent the background and observation error covariance matrices, respectively. The gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$, is computed via automatic differentiation. Computing the gradient of the 4DVar cost function projects observational information onto the model state space, quantifying the necessary adjustments to the background field. Consequently, this process serves as a unified information interface for heterogeneous observations, obviating the need to design specialized encoders for distinct data types.
By combining the gradient $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ with the background field $\mathbf{x}^{b}(t_{0})$, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. Among them, our DA task involves sequentially employing DA models for different observational data. For example, in this study, XiChen first assimilates GDAS prepbufr data, then assimilates MHS, ASCAT, AMSU-A, and SATWND observations. This configuration enhances XiChen’s flexibility. The assimilation of new observations in the future merely requires fine-tuning the respective observation operators and DA model components, eliminating the need to retrain the entire DA system. For details of the cascade DA framework, please refer to the Supplementary Text and Figure S1 (b).
4.4 Pre-training for medium-range global weather forecasting
@@
We aim to design a foundation model that can be pre-trained on the reanalysis data and then be fine-tuned to solve various downstream weather tasks to build a global weather observation-to-forecast ML system. As illustrated in Figure 1(c), during the pre-training process, the foundation model takes the reanalysis field $\mathbf{x}(t)$ as the input and the lead time $\Delta t$ as the condition and outputs the predicted state $\hat{\mathbf{x}}(t+\Delta t)$. In this study, we sample $\Delta t$ from the set $\{1,3,6,12,24\}$ hours. This process enables the foundation model to effectively capture the relationships among various weather variables while accounting for both the dynamical and thermodynamic characteristics of the atmosphere. To generate predictions for any time increments, the predictions from the learned forecasting model can be combined using the hierarchical temporal aggregation methodology proposed by Pangu-Weather [bi2023accurate]. Thus, the foundation model is pre-trained to minimize the following loss:
@@
$\tiny{\mathcal{L}(\theta)=\mathbb{E}\left[\frac{1}{VHW}\sum^{V}_{v=1}\sum^{H}_{i=1}\sum^{W}_{j=1}\omega(v)L(i)\|\Delta\mathbf{x}_{v,i,j}(t+\Delta t)\|_{1}\right],}$ (2)
@@
where $\omega(v)$ denotes the pressure-weighting coefficient of variable $v$ [lam2023learning], $\Delta\mathbf{x}(t+\Delta t)=\mathcal{M}_{t\rightarrow t+\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\Delta t)$ is the difference between the predicted state and the ground truth at time $t+\Delta t$, $\|\|_{1}$ represents the $L_{1}$-norm, and $L(i)$ is the latitude-weighting factor [rasp2020weatherbench]:
@@
where $\text{lat}(i)$ is the latitude of the $i$th row of the grid. This term is commonly used in the training process of previous ML-based weather forecasting models [kurth2023fourcastnet, bi2023accurate, lam2023learning]. This pre-training process costs about 45 hours to run on 8 A100-40GB GPUs.
@@
where $\tau$ denotes the roll-out step and $\Delta\mathbf{x}_{v,i,j}(t+\tau\Delta t)=\mathcal{M}_{t\rightarrow t+\tau\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\tau\Delta t)$.
In practice, we first fine-tune the pre-trained forecasting model with roll-out steps from $T=2$ to $T=4$ to develop the XiChen-Short model. The weights obtained from the XiChen-Short model are then used to initialize the XiChen-Medium model, which is subsequently fine-tuned to achieve optimal forecasting performance for periods of 5 to 10 roll-out steps. We use the same sampled value of $\Delta t$ for all $T$ steps. Our forecasting model is similar to FuXi [fudanchen2023fuxi] in its forecasting process. Since the analysis field errors provided by the XiChen DA model are comparable to those of XiChen-Short’s 2-day forecasts initialized with ERA5, the 10-day medium-range global weather forecasts using initial fields estimated by the XiChen DA model are generated by XiChen-Short for the first 3 days and by XiChen-Medium for the remaining 7 days. The fine-tuning process of the XiChen forecasting model costs about 94 hours to run on 8 A100-40GB GPUs. For additional information on the model training configuration, please refer to the Supplementary Text.
@@ 4.5 Fine-tuning for the observation operator
To compare the atmospheric state with satellite observations, the observation operator is essential [liang2023machine]. The simulated radiance is computed at each observation location using a profile of atmospheric variables. As illustrated in Figure 1(d), we fine-tune the pre-trained model by taking the satellite’s attitude parameters and scan position as conditions to construct an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. The loss function used to train $\hat{\mathcal{H}}^{s}$ is as follows:
@@ 4.6 Fine-tuning for data assimilation
As illustrated in Figure 1(e), we fine-tune the pre-trained model to function as a DA model by incorporating the gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$. This approach follows the derivation presented in 4DVarFormer [wang2024accurate]. However, unlike 4DVarFormer, XiChen outputs the representation of analysis increments in latent space rather than the analysis increments themselves. This representation is added to the embedding of the background field $\mathbf{x}^{b}(t_{0})$ and subsequently decoded by the HNO blocks to generate the analysis field.
@@
To circumvent the temporal latency inherent in ML-based 4DVar methods [xiao2024fengwu4dvar, wang2024accurate] when deploying XiChen as a self-contained forecasting system for acquiring medium-range prognostic results, we have architected a dual assimilation framework (see the Supplementary Text and Figure S4 for details). This dual DA framework facilitates the real-time acquisition of medium-range global weather forecasting results without necessitating the temporal delay associated with the availability of all observations within the complete 12-hour DAW. Figures S9 and S10 illustrate the impact of varying assimilation windows on both DA cycle performance and 10-day medium-range forecast performance.